# Mitigating Text Toxicity with Counterfactual Generation

### v1

@@ Abstract
Toxicity mitigation consists in rephrasing text in order to remove offensive or harmful meaning. Neural natural language processing (NLP) models have been widely used to target and mitigate textual toxicity. However, existing methods fail to detoxify text while preserving the initial non-toxic meaning at the same time. In this work, we propose to apply counterfactual generation methods from the eXplainable AI (XAI) field to target and mitigate textual toxicity. In particular, we perform text detoxification by applying local feature importance and counterfactual generation methods to a toxicity classifier distinguishing between toxic and non-toxic texts. We carry out text detoxification through counterfactual generation on three datasets and compare our approach to three competitors. Automatic and human evaluations show that recently developed NLP counterfactual generators can mitigate toxicity accurately while better preserving the meaning of the initial text as compared to classical detoxification methods. Finally, we take a step back from using automated detoxification tools, and discuss how to manage the polysemous nature of toxicity and the risk of malicious use of detoxification tools. This work is the first to bridge the gap between counterfactual generation and text detoxification and paves the way towards more practical application of XAI methods.
@@ 1 Introduction
Online textual toxicity can be considered as rude, aggressive and degrading attitudes exhibited on online platforms, ranging from harmful to hateful speech. Hate speech is defined as aggressive or offensive language against a specific group of people who share common characteristics, such as religion, race, gender, sexual orientation, sex or political affiliation Castaño-Pulgarín et al. (2021). Such toxic content has multiplied on the Internet in recent years Thomas et al. (2021), raising concerns about its multi-faceted negative impact, such as the potential to threaten victims’ psychological and physical well-being Walther (2022) or to be used as a medium for criminal actions Rapp (2021).
@@
Toxic text data can also have a negative impact when used to train large language models (LLMs): recent advances in natural language processing (NLP) and the development of LLMs such as GPT-3 Brown et al. (2020), or LaMDA Thoppilan et al. (2022) have been made possible by utilizing vast quantities of textual data available on the Internet. These models have demonstrated a high capacity to generate plausible text, while raising several concerns about harmful content generation Bender et al. (2021), and bias amplification Mehrabi et al. (2022); Gallegos et al. (2023) coming from the training texts. Thus, LLMs’ ability to generate toxic content may contribute to the rapid spread of such content online, as more and more content is synthetically generated by chatbots Martínez et al. (2023).
@@
To overcome the rapid development of online toxic content and curb its societal impact, automatic toxicity processing methods have been developed to detect and process harmful content on online communities and digital media platforms Fortuna and Nunes (2018). In particular, text detoxification (or toxicity mitigation) aims to rewrite toxic text in order to remove (or mitigate) toxicity while preserving the initial non-toxic meaning and maintaining plausibility. Several methods based on neural NLP models have been developed to perform text detoxification Hallinan et al. (2023); Dale et al. (2021); Laugier et al. (2021) by generating text under constraint, or by detecting toxic content and modifying it. If these methods succeed in significantly lowering textual toxicity, they generally fail to preserve the initial non-toxic content. Besides, automatic toxicity processing tools raise major ethical questions regarding the risks related to their robustness and the role of humans involved.
@@
In this paper we propose to address toxicity targeting and mitigation by applying eXplainable AI (XAI) methods, more precisely Local Feature Importance (LFI) and counterfactual example generation Molnar (2020). The former aims at detecting important input features to explain a model prediction. Counterfactual example generation (see Guidotti (2022) for a survey) instead explains a model’s prediction by identifying the minimal changes that enable flipping the outcome of a classifier.
@@
We show that LFI and counterfactual generation can be applied to a toxicity classifier respectively to target toxicity and to perform toxicity mitigation.
We propose CF-Detox, a toxicity mitigation method based on a recently developed counterfactual example generator: TIGTEC Bhan et al. (2023a).
We conduct both automatic and human experiments to show that CF-Detox reaches competitive performance in text detoxification.
We discuss risks and opportunities related to automatic toxicity detection and mitigation tools and define recommendations.
As an illustration, Figure 1 shows an initial toxic text and its detoxified versions respectively obtained from our proposed method, MaRCo Hallinan et al. (2023) and CondBERT Dale et al. (2021). Counterfactual detoxification leads here to more sparse, plausible and context preserving text as compared to the other methods.
The paper is organized as follows: in Section Background and Related Work we first recall basic elements about toxicity mitigation and summarize the desired characteristics of detoxification methods. The second section shows how common XAI methods can be used to (1) target toxicity and (2) generate plausible and content-preserving detoxified texts. Experimental results discussed in the next section highlight that text detoxification through counterfactual generation achieves competitive results in terms of toxicity mitigation, content preservation and plausibility, as compared to state-of-the-art competitors. It also compares the ability of different LFI methods to target toxic content. Experimental evaluation is performed, both automatically and with a human-grounded protocol. Since the use of automatic toxicity processing raises several critical concerns, we finally discuss in the last section risks and opportunities around the use of toxicity mitigation methods. As a result, we discuss how to take full account of the polysemous nature of toxicity, manage the risk of malicious use of detoxification tools and favor human-in-the-loop processes.
@@ 2 Background and Related Work
In this section we recall the context of automatic text detoxification, whose objective is to remove toxicity while preserving the initial non-toxic content. We present the task of text detoxification and discuss existing methods that aim to detoxify text using neural NLP models. Finally, we introduce the XAI principles used in the next section to perform toxicity detection and mitigation.
@@ 2.1 Automatic toxicity processing background
In the following, we use the terms text detoxification and toxicity mitigation interchangeably, as their definitions depend on that of toxicity.
@@ 2.1.1 Definition and objective
Textual toxicity can be defined in multiple ways Fortuna and Nunes (2018) and can take various forms, such as rude, offensive or hateful speech, potentially causing online harm to isolated people or minority groups Thomas et al. (2021). Automatic toxicity processing can essentially take two forms: detection and mitigation. Toxicity detection can be either based on prior knowledge (vocabulary, regex) or obtained from a fine tuned toxicity classifier mapping an input text representation space to an output space to distinguish toxic and non-toxic texts Cook et al. (2023). Training language models to classify toxic and non-toxic texts is difficult, as it requires access to datasets labeled based on an implicit definition of toxicity derived from human annotators Jigsaw (2018). Toxicity mitigation consists in rewriting a toxic text while preserving the non-toxic meaning. This task is even more difficult because it requires disentangling toxic and non-toxic meanings to plausibly modify the former while preserving the latter.
2.1.2 Expected characteristics of text detoxification
Several desirable properties have been proposed to assess automatically text detoxification. We organize them into three categories. When ground truth detoxified text is unavailable, we use the previously introduced classifier as an oracle to evaluate the toxicity level of the supposedly detoxified text. Accuracy (ACC) assesses the extent to which the generated texts are accurately detoxified with respect to . Accuracy can be measured by computing either the rate of successful changes or the average toxicity logit of the generated texts using .
Proximity or content preservation (CP) evaluates how close two texts are. Textual similarity can be defined in two different ways. The first one consists of evaluating textual proximity based on word sequence co-occurrences, with metrics such as self-BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Denkowski and Lavie (2011) or the Levenshtein distance. The second way of evaluating textual proximity is to measure semantic similarity from word-level embeddings Pennington et al. (2014) or sentence-level embeddings Zhang et al. (2019); Wieting et al. (2019).
@@ 2.2 Toxicity mitigation with neural NLP models
This section presents existing methods leveraging neural NLP models to classify and generate plausible text for text detoxification. Two categories are distinguished: Text Style Transfer (TST) and Masking and Reconstructing (M&R).
@@ 2.2.1 Text Style Transfer
TST (see Hu et al. (2022) for a survey) aims to alter the stylistic attributes of an initial text while preserving its content that is unrelated to the target style. Text detoxification can be achieved through TST, where the initial style is characterized by the presence of toxicity, and the target style is defined by its absence. Style transfer is usually performed by generating text with neural NLP decoders guided by an NLP toxicity classifier. In general, TST methods vary based on the language model used for text generation and the method of text generation steering.
A first TST approach Nogueira dos Santos et al. (2018) uses an encoder-decoder architecture based on recurrent neural networks (RNN) to generate non-toxic text, using a toxicity convolutional neural network (CNN) classifier to steer the style transfer. Another method Laugier et al. (2021) fine-tunes a text-to-text T5 model Raffel et al. (2019) by using a denoising and cyclic auto-encoder loss. Finally, ParaGeDi Dale et al. (2021) uses a pre-trained T5-based paraphraser model and a class-conditionned language model to steer the text generation.
TST methods generally detoxify text accurately but struggle to preserve its non-toxic meaning Hallinan et al. (2023).
@@ 2.2.2 Masking and Reconstructing
Toxicity mitigation can be sequentially done by (1) targeting toxic content, (2) masking it, and (3) modifying it. Masking and Reconstructing (M&R) approaches generally enable performing text detoxification while preserving the non-toxic meaning. Once the toxic content is targeted, mask infilling is usually performed with a neural NLP encoder. In general, M&R methods differ in the way they target toxic content, and the neural NLP model used to perform mask infilling.
A first M&R approach Tran et al. (2020) performs text detoxification by retrieving potential harmful Part-Of-Speech (POS) based on a predefined vocabulary of toxic words, generating non-offensive POS substitution candidates, and editing the initial text through mask infilling with a RoBERTa NLP encoder for unacceptable candidates. CondBERT Dale et al. (2021) identifies tokens to be masked using a logistic bag-of-words classifier and performs mask infilling using a BERT NLP encoder. Finally, the most recent M&R method called MaRCo Hallinan et al. (2023) detects POS that could convey toxic meaning by comparing likelihoods from two BART NLP encoder-decoder respectively fine-tuned on toxic and non-toxic content. The targeted potential toxic content is then replaced by non-toxic content by mixing token probabilities from these two encoder-decoders and a third neutral model.
On average, M&R yields to better results than TST in terms of content preservation, and performs equally regarding toxicity mitigation Hallinan et al. (2023).
@@ 2.3 XAI for NLP
In the following, we consider the neural NLP toxicity classifier introduced in the previous section, and a text represented as a sequence of tokens with . can either be a binary space that distinguishes toxic and non-toxic texts or a multi-class space that categorizes several levels of toxicity.
@@ 2.3.1 Local Feature Importance
A Local Feature Importance (LFI) function explains a prediction by a vector where is the contribution of the th token to the prediction. The higher the contribution, the more important the token to explain the prediction of the classifier . Three types of LFI methods can be distinguished: perturbation-based such as KernelSHAP Lundberg and Lee (2017), gradient-based such as Integrated gradients Sundararajan et al. (2017) and attention-based such as self-attention in case of a Transformer classifier Bhan et al. (2023b). Perturbation-based LFI perturbs and resamples feature values to compute feature importance, whereas gradient-based LFI is based on the classifier’s backpropagated gradient activity.
@@ 2.3.2 Counterfactual explanations
Counterfactual explanations emphasize what should be different in an input instance to change the outcome of a classifier Ali et al. (2023). Counterfactual examples provide contrastive explanations by simulating alternative causes to assess if a specific event (here the predicted class) still happens or not Miller (2019). The counterfactual example generation can be formalized as a constrained optimization problem. For a given classifier and an instance of interest , a counterfactual example must be close to but predicted differently. It is defined as:
where is a cost function that aggregates several expected counterfactual characteristics, such as the textual distance. The counterfactual explanation is then the difference between the generated counterfactual example and the initial data point, . Many desirable characteristics for counterfactual explanations have been proposed Guidotti (2022); de Oliveira and Martens (2021), such as sparsity, defined as the norm of , plausibility to make sure that the counterfactual example is not out-of-distribution Laugel et al. (2019) and actionability Poyiadzi et al. (2020). In the following section, we present a few methods for generating textual counterfactuals by comparing and relating them with the toxicity mitigation task.
@@ 3 When XAI meets text detoxification
This section describes how XAI methods can be used to tackle text detoxification. We show that LFI methods can foster toxic content targeting and we illustrate how to apply counterfactual generation methods for performing text detoxification.
@@ 3.1 Targeting toxic POS with Local Feature Importance
Toxic POS targeting involves identifying the elements in a toxic text that induce its toxicity. While toxicity can be easily defined in part by a predefined lexical field, it can also take more complex forms, such as sarcasm, synonyms, or associations that are difficult to detect automatically.
Let be a toxicity classifier and is a toxic text with . Applying LFI to highlights important tokens that explain why has been classified as toxic by . This way, toxic POS detection can be performed by applying LFI methods to on texts classified as toxic.
@@
Toxic POS detection with LFI does not require the definition of a predefined toxicity vocabulary and is only based on a model that is trained to discriminate between toxic and non-toxic texts. Then, toxicity is detected in a data driven fashion based on a fine-tuned neural NLP model. The use of LFI methods applied to makes it possible to detect complex forms of toxicity, as recent neural NLP models such as BERT take into account the context to make their predictions. Toxic POS detection through LFI depends on the ability of the classifier to accurately discriminate between toxic and non-toxic texts. A toxicity classifier with a low accuracy might misclassify toxic and non-toxic texts, leading to non-toxic important tokens being highlighted with LFI explanations.
@@
Figure 2 illustrates the principle of toxic POS targeting with LFI applied to a toxicity classifier, where the tokens "f**k" and "b***h" are assessed as important to predict that the first text is toxic, whereas the tokens "ugliest" and "b***h" are highlighted for the second text. In the next section, we experimentally study the relevance of this approach by comparing perturbation-based, attention-based and gradient-based LFI methods.
3.2 Toxicity mitigation with counterfactual generation
We propose to perform toxicity mitigation through counterfactual generation with respect to the toxicity classifier. We postulate that, following the notations of Equation 1, toxicity mitigation can be performed by setting as a toxic instance of interest with . Therefore, the objective is to find that minimizes function such that . The lack of ground truth detoxified texts is then overcome through the use of as an oracle to guide text detoxification while keeping the non-toxic content. This way, text detoxification through counterfactual generation consists in detecting texts classified as toxic by and generating their related detoxified counterfactual examples.
Among the numerous expected attributes of counterfactual examples, we highlight those that are common with the expected characteristics of text mitigation introduced in the previous section:
Proximity is identical to content preservation (CP) that can be measured in NLP with either sparsity (number of token changes between the initial text and its related counterfactual example) or semantic similarity (computed from sentence embedding similarity).
Plausibility that can be treated in NLP as the linguistic fluency and measured with the perplexity (PPL).

### v2

@@ Abstract
Toxicity mitigation consists in rephrasing text in order to remove offensive or harmful meaning. Neural natural language processing (NLP) models have been widely used to target and mitigate textual toxicity. However, existing methods fail to detoxify text while preserving the initial non-toxic meaning at the same time. In this work, we propose to apply eXplainable AI (XAI) methods to both target and mitigate textual toxicity. We propose CF-Detox to perform text detoxification by applying local feature importance, counterfactual example generation and counterfactual feature importance methods to a toxicity classifier distinguishing between toxic and non-toxic texts. We carry out text detoxification through counterfactual generation on three datasets and compare our approach to three competitors. Automatic and human evaluations show that recently developed NLP counterfactual generators can mitigate toxicity accurately while better preserving the non-toxic meaning of the initial text as compared to classical detoxification methods. Finally, we take a step back and discuss ethical considerations on how to manage the diversity of toxicity perception and the risk of malicious use of detoxification tools. This work is the first to bridge the gap between counterfactual generation and text detoxification and paves the way towards more practical applications of XAI methods.
@@ 1 Introduction
Online textual toxicity can be considered as rude, aggressive and degrading attitudes exhibited on online platforms, ranging from harmful to hateful speech. Hateful speech is defined as aggressive or offensive language against a specific group of people who share common characteristics, such as religion, race, gender, sexual orientation, sex or political affiliation Castaño-Pulgarín et al. (2021). Such toxic content has proliferated on the Internet in the recent years Thomas et al. (2021), raising concerns about its multi-faceted negative impact, such as the potential to threaten the psychological and physical well-being of victims Walther (2022) or to be used as a medium for criminal actions Rapp (2021).
@@
Toxic text data can also have a negative impact when used to train large language models (LLMs). Recent advances in natural language processing (NLP) and the development of LLMs such as GPT-3 Brown et al. (2020) have been made possible by utilizing vast quantities of textual data available on the Internet. These models have demonstrated a high capacity to generate plausible text, while raising several risks about harmful content generation Bender et al. (2021) and bias amplification Gallegos et al. (2023) coming from the training texts. Thus, by generating toxic content, LLMs may contribute to the rapid spread of online toxicity, as as text is increasingly generated synthetically by chatbots Martínez et al. (2023).
@@
To cope with the rapid development of online toxic content and curb its societal impact, automatic toxicity processing methods have been developed to detect and process hateful content in online communities and digital media platforms Fortuna and Nunes (2018). For example, such tools can be used to clean up language model training datasets, or to suggest interventions on toxic and hateful messages to the human-labor moderating online platforms. In particular, text detoxification (or toxicity mitigation) aims to rewrite toxic text in order to remove (or to mitigate) toxicity while preserving the initial non-toxic meaning and maintaining grammatical plausibility. Recent methods based on neural NLP models have been developed to perform text detoxification Hallinan et al. (2023); Dale et al. (2021); Laugier et al. (2021) by either generating text under constraint, or by detecting toxic content and modifying it. While these methods succeed in drastically lowering textual toxicity, they generally fail to preserve the initial non-toxic content. Besides, automatic online toxicity processing tools raise major ethical questions regarding the risks related to their robustness and their relationship with online human-labor.
@@
In this paper we propose to address toxicity targeting and mitigation by applying eXplainable AI (XAI) methods Molnar (2020): more precisely Local Feature Importance (LFI) Barredo Arrieta et al. (2020), counterfactual (CF) generators (see Guidotti (2022) for a survey) and Counterfactual Feature Importance (CFI) Bhan et al. (2023a). The former aims at detecting important input features to explain a model prediction. Instead, CF example generation explains a model prediction by identifying the minimal changes that enable flipping the outcome of a classifier. CFI measures token modification importance to explain label flipping related to CF explanations.
@@
We show that LFI and CF generation can be applied to a toxicity classifier to target toxicity and to perform toxicity mitigation respectively,
We propose CF-Detox, a detoxification method based on a recently developed CF generator: TIGTEC Bhan et al. (2023b) and validated by both automatic and human experiments,
We illustrate how to apply CFI methods to make detoxified texts even more content preserving,
We discuss risks and opportunities related to automatic toxicity processing tools and define recommendations.
For illustrative purposes, Figure 1 shows an initial toxic text and its detoxified versions obtained from our proposed method, MaRCo Hallinan et al. (2023) and CondBERT Dale et al. (2021) respectively. Here, counterfactual detoxification leads to more plausible, sparse and context preserving text as compared to the other methods.
The paper is organized as follows, in Section 2 we first recall basic principles about automatic toxicity processing and XAI. In Section 3 we propose methods to use common XAI approaches to (1) target toxicity and (2) generate plausible and content-preserving detoxified texts. Experimental results discussed in Section 4 highlight that text detoxification through CF generation achieves competitive results in terms of toxicity mitigation, content preservation and plausibility, as compared to state-of-the-art competitors. This section also compares the ability of different LFI methods to target toxic content. Experimental evaluation is performed, both automatically and with a human-grounded protocol. After discussion and conclusion, we finally discuss in Section 7 risks and opportunities around the use of toxicity mitigation methods. As a result, we discuss how to manage the diversity of toxicity perception and the risk of malicious use of detoxification tools and favor human-in-the-loop processes.
@@ 2 Background and Related Work
This section recalls the task of automatic toxicity detection and mitigation. We presents existing methods aiming to detoxify text using neural NLP models. We then the XAI principles used in the next section to perform toxicity detection and mitigation. In the following, we employ text detoxification and toxicity mitigation interchangeably.
@@ 2.1.1 Definition and objective
Textual toxicity can be defined in multiple ways Fortuna and Nunes (2018) and can take various forms, such as rude, offensive or hateful speech, potentially causing online harm to isolated people or minority groups Thomas et al. (2021). Automatic toxicity processing can essentially take two forms: detection and mitigation. Toxicity detection can be either based on prior knowledge (vocabulary, regex) or obtained from a fine tuned toxicity classifier mapping an input text representation space to an output space to distinguish toxic and non-toxic texts Cook et al. (2023). Training language models to perform this classification is difficult, as it requires access to datasets labeled based on an idiosyncratic definition of toxicity derived from human annotators Jigsaw (2018). Toxicity mitigation consists in rewriting a toxic text while preserving the non-toxic meaning. This task is even more difficult because it requires disentangling toxic and non-toxic meanings to modify the former plausibly while preserving the latter.
2.1.2 Evaluation criteria of text detoxification
Numerous desirable properties have been proposed to assess automatically text detoxification. We organize them into three categories. When ground truth detoxified texts are unavailable, we use the previously introduced classifier as an oracle to evaluate the toxicity level of the supposedly detoxified text. Accuracy (ACC) measures the extent to which the generated texts are accurately detoxified with respect to . Accuracy can be measured by computing either the rate of successful changes or the average toxicity score of the generated texts using . Proximity or content preservation (CP) evaluates how close two texts are. Textual similarity can be defined in two ways. The first one consists in evaluating textual proximity based on word sequence co-occurrences, with metrics such as self-BLEU Papineni et al. (2002) or the Levenshtein distance. The second consists in measuring semantic similarity from word-level embeddings Pennington et al. (2014) or sentence-level embeddings Zhang et al. (2019). Finally, text plausibility is automatically evaluated by the perplexity (PPL) score usually obtained from generative language models such as GPT-2 Radford et al. (2018).
@@ 2.2 Toxicity Mitigation with Neural NLP Models
This section describes two categories of methods leveraging neural NLP models to perform detoxification, namely Text Style Transfer (TST) and Masking and Reconstructing (M&R).
@@ 2.2.1 Text Style Transfer
TST (see Hu et al. (2022) for a survey) aims to alter the stylistic attributes of an initial text while preserving its content that is unrelated to the target style. Text detoxification can be achieved through TST, where the initial style is characterized by the presence of toxicity, and the target style is defined by its absence. TST is usually performed by generating text with neural NLP decoders guided by an NLP toxicity classifier. In general, TST methods vary based on the language model used for autoregressive text generation and the heuristic of text generation steering.
A first TST approach Nogueira dos Santos et al. (2018) uses an encoder-decoder architecture based on recurrent neural networks to generate non-toxic text, using a toxicity convolutional neural network classifier to steer the style transfer. Another method Laugier et al. (2021) fine-tunes a text-to-text T5 model Raffel et al. (2019) by using a denoising and cyclic auto-encoder loss. ParaGeDi Dale et al. (2021) uses a pre-trained T5-based paraphraser model and a class-conditioned language model to steer the text generation. TST methods generally detoxify text accurately but struggle to preserve its non-toxic meaning Hallinan et al. (2023).
@@ 2.2.2 Masking and Reconstructing
Masking and Reconstructing (M&R) approaches perform toxicity mitigation by sequentially (1) targeting toxic content, (2) masking it, and (3) modifying it. Once the toxic content is targeted, mask infilling is usually performed with a neural NLP encoder. In general, M&R methods differ in the way they target toxic content, and the neural NLP model used to perform mask infilling.
A first M&R approach Tran et al. (2020) detoxifies text by retrieving potential harmful Part-Of-Speech (POS) based on a predefined toxicity vocabulary, generating non-offensive POS substitution candidates, and editing the initial text through mask infilling with a RoBERTa encoder for unacceptable candidates. CondBERT Dale et al. (2021) identifies tokens to be masked using a logistic bag-of-words classifier and performs mask infilling using a BERT encoder. The most recent M&R method called MaRCo Hallinan et al. (2023) detects POS that could convey toxic meaning by comparing likelihoods from two BART encoder-decoders respectively fine-tuned on toxic and non-toxic content. The targeted potential toxic content is then replaced by non-toxic content by mixing token probabilities from these two encoder-decoders and a third neutral model. On average, M&R yields to better results than TST in terms of content preservation, and performs equally regarding toxicity mitigation Hallinan et al. (2023).
@@ 2.3 XAI for NLP
In the following, we consider the neural NLP toxicity classifier introduced in the previous section, and a text represented as a sequence of tokens of length with . can either be a binary space that distinguishes toxic and non-toxic texts or a multi-class space that categorizes several levels of toxicity.
@@ 2.3.1 Local Feature Importance
A Local Feature Importance (LFI) function explains a prediction by a vector where is the contribution of the th token to the prediction. The higher the contribution, the more important the token is to explain the prediction of the classifier . Three types of LFI methods can be distinguished: perturbation-based such as KernelSHAP Lundberg and Lee (2017), gradient-based such as Integrated gradients Sundararajan et al. (2017) and attention-based such as self-attention in case of a Transformer classifier Bhan et al. (2023c).
@@ 2.3.2 Counterfactual explanations
Counterfactual (CF) explanations emphasize what should be different in an input instance to change the associated outcome of a classifier Ali et al. (2023). CF examples provide contrastive explanations by simulating alternative instances to assess whether a specific event (in our case the predicted class) still occurs or not Miller (2019). The CF example generation can be formalized as a constrained optimization problem. For a given classifier and an instance of interest , a CF example must be close to but predicted differently. It is defined as:
where is a cost function that aggregates several expected CF characteristics, such as the textual distance. The CF explanation is then the difference between the generated CF example and the initial data point, . Many desirable characteristics for CF explanations have been proposed Guidotti (2022), such as sparsity or plausibility to make sure that the CF example is not out-of-distribution Laugel et al. (2019). In Section 3, we present several methods to generate textual CFs by comparing and relating them with the toxicity mitigation task.
2.3.3 Counterfactual Feature Importance
A Counterfactual Feature Importance (CFI) operator computes token substitution importance between an instance of interest and its related CF explanation Bhan et al. (2023a). In the same way as common LFI, is the contribution of the -th token modification. CFI highlights the most important token modifications, making non-sparse CF explanations easier to read and more intelligible. A simple but effective way to compute CFI is to apply Integrated gradients to explain the difference between and , by setting the baseline to the initial instance of interest .
@@ 3 When XAI meets text detoxification
In this section we show how LFI methods can foster toxic content targeting and discuss how to apply CF generation methods for performing text detoxification. We also propose a new method based on CFI to make detoxified texts even more content preserving.
@@ 3.1 Targeting Toxic POS with Local Feature Importance
Toxic POS targeting consists in identifying the elements in a toxic text that induce its toxicity. While toxicity can be easily defined partially by a predefined lexical field, it can also take on more complex forms implying specific cultural references or sarcasm that are difficult to detect automatically. Let be a toxicity classifier and is a toxic text with .
@@
We propose to apply LFI to to highlight important tokens that explain why has been classified as toxic by , enabling toxic POS detection. This way, toxic POS detection with LFI does not require the definition of a predefined toxicity vocabulary and is only based on a model that is trained to discriminate between toxic and non-toxic texts. Then, toxicity is detected in a data-driven manner based on a fine-tuned neural NLP model. Our proposition to apply LFI methods to makes it possible to detect complex forms of toxicity, as recent neural NLP models such as BERT can encode high level linguistic forms to make their predictions. Toxic POS detection through LFI depends on the ability of the classifier to accurately discriminate between toxic and non-toxic texts. Consequently, a toxicity classifier with a low accuracy might misclassify toxic and non-toxic texts, leading to unreliable explanations.
@@
The top part of Figure 2 (toxicity targeting) shows an example of toxic POS targeting with LFI applied to a toxicity classifier: the token "f**k" is assessed as important to predict toxicity. In Section 4, we experimentally study the relevance of this approach by comparing perturbation-based, attention-based and gradient-based LFI methods.
3.2 Epistemic Similarities Between Text Detoxification and Counterfactual Generation
In the following, we disclose that toxicity mitigation and CF generation share common features in terms of objective, evaluation criteria and categories of approaches.
3.2.1 Similarity of objectives and evaluation criteria.
Firstly, the aim of both toxicity mitigation and CF generation is to generate text close to an initial instance, while reaching a target state. For detoxification purpose, the target state is defined as non-toxicity, whereas the target state is defined by the classifier label for CF generation. Secondly, as discussed in Section 2, toxicity mitigation methods share with CF generators several evaluation criteria, such as accuracy, proximity and plausibility. These evaluation criteria are measured with the same metrics, such as sparsity, embedding-based semantic similarity, BLEU score or perplexity measures.
3.2.2 Similarity of constitutive categories of methods.
Thirdly, methods detoxifying text and generating CFs can be grouped into two similar categories. As mentioned in Bhan et al. (2023b), textual CF generators can be of two types: Text editing heuristics and Counterfactual generation with large language models.